A number consist of two digit whose sum is 5. When the digits are reversed, tne number becomes greater by 9.Find the number:
Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(75887)

A number consist of two digit whose sum is 5. When the digits are reversed, tne number becomes greater by 9.Find the number:
-----
Let the number be 10t+u where t is the tens and u is the units digit.
----
Equations:
t + u = 5
10u+t = 10t+u +9
---------------------
Simplify :
t + u = 5
9u - 9t = 9
----
t + u = 5
-t+u = 9
---
Add:
2u = 14
u = 7
---
But, if u = 7, then t= -2
and that makes no sense
for a 2-digit number.
=============================
Cheers,
Stan H.

Answer by Alan3354(69443)

A number consist of two digit whose sum is 5. When the digits are reversed, tne number becomes greater by 9.Find the number:
--------------
It has to be 14 or 23
It's 23
--------------
The difference when you reverse digits is always a multiple of 9.
1*9
2*9
3*9
4*9 etc
The multiplier of 9 is the difference between the digits.
eg, for 37, the difference is 4, so 4*9 will be the change when the digits are reversed.
-------------
For your problem, the multiplier of 9 is 1, so the difference between the digits is 1.
That makes the problem simpler:
---------------------------------
The sum of the digits is 5
The difference of the digits is 1
u + t = 5
u - t = 1
------------ Add
2u = 6
u = 3
t = 2
--> 23
-------------------
The long way:
10t + u = 10u + t - 9
t + u = 5
----------
9t - 9u = -9
t - u = -1
t + u = 5
2t = 4
t = 2
u = 3
--> 23